# Singapo: Single Image Controlled Generation of Articulated Parts in Objects

### v2

@@ 1 Introduction
Figure 1: Our work proposes to generate 3D articulated objects from a single image oberving the object in the resting state from a random view. Left: We design a pipeline to synthesize articulated assets progressively from coarse to fine details in a modular way. Right: Our method generates objects with varying part geometry and motion to account for the ambiguity in the input image.
@@ 2 Related Work
Conditional generation on 3D structures. Enabling control over the generation process is crucial for practical applications. This becomes more challenging for structured data generation, due to the need for fine-grained control and parsing the conditioning into structured components. Part-based object generation has been studied in settings driven by a partial shape to complete the object part-by-part or guided by an image (Wu et al., 2020; Niu et al., 2018). These methods only work with specific categories of objects that share similar structures. The recent advances in large-language models and generative models promote conditional generation from 2D (Yang et al., 2021; Rombach et al., 2022; Zhang et al., 2023a; Blattmann et al., 2023) to 3D with structures (Liu et al., 2024a). Many works on 3D scene generation have explored the control of object arrangements via text for room and object specification (Zhang et al., 2023b; Lin & Mu, 2024; Aguina-Kang et al., 2024; Tam et al., 2024), 2D/3D layout for spatial constraint (Bahmani et al., 2023; Eldesokey & Wonka, 2024; Feng et al., 2024; Po & Wetzstein, 2024; Maillard et al., 2024; Yang et al., 2024), and scene graphs as spatial or semantic priors (Lin & Mu, 2024; Zhai et al., 2024; Gao et al., 2024). Chen et al. (2024a) generate objects from an image by decomposing the subscene into objects to generate separately. Alam & Ahmed (2024) propose to generate parametric CAD command sequences conditioned on an image. Our work follows the trend of conditional generation on structured data. To the best of our knowledge, we are the first to explore articulated object generation from a single image.
@@ 3 Methods
Figure 2: Our method takes an object image as input and generates attributes of articulated parts, which are used to assemble the object via part mesh retrieval. We design a diffusion-based model for part generation, which is guided by a part connectivity graph and DINO patch features of the input image. Our denoising network is built on layers of attention blocks. The graph constraint is injected into the graph relation module by converting to an adjacency matrix as the attention mask. The image features act as the keys and values in the cross attention to condition the part arrangement.
@@ 3.2 Data Representation
We assume objects to be normalized inside a cube of size $2$ centered at the origin and all the parameters are defined in the world coordinate system. We describe each articulated part with multiple attributes $\mathbf{p}_{i}=\{\mathbf{b}_{i},l_{i},t_{i},\mathbf{a}_{i},\mathbf{r}_{i}\}$ at abstract level, including the part bounding box $\mathbf{b}_{i}\in\mathbb{R}^{6}$, semantic label $l_{i}$, articulation type $t_{i}$, location and direction of joint axis $\mathbf{a}_{i}\in\mathbb{R}^{6}$, and motion range $\mathbf{t}_{i}\in\mathbb{R}^{2}$. The articulation type considered in this work includes fixed, revolute, prismatic, continuous and screw joints. The semantic label includes base, door, drawer, handle, knob, tray. The semantic label will be used to retrieve the part geometry from the dataset. Spatially, the parts are organized into a connectivity graph, where each node represents a part, and the edge direction indicates the parent-child relationship. This spatial relationship also defines the kinematic constraints between parts. During training, the graph is represented as an adjacency matrix $A\in\{0,1\}^{N\times N}$, where $A_{ij}=1$ if part $i$ is in connection with part $j$. The matrix $A$ will serve as an attention mask to guide the part generation process. To handle a different number of parts in the denoising diffusion process, we pad the parts to a maximum number $32$ with all part attributes represented in $6$ dimensions.
@@ 3.3.2 Model architecture
Classifier-free training for conditioning. Our denoising network takes image $\mathcal{I}$ as the main control and graph $G$ as auxiliary guidance. Both of them are critical as the image describes the spatial arrangement of parts and the graph determines the ordering of the parts to be generated so that the parts are organized correctly when articulating. Inherent from the base model, our network can also take object category $c$ as an additional condition, which is fused with the timestep embeddings via an adaptive norm layer (Xu et al., 2019) prior to attentions in each block. We empirically find that the original design of CAGE that takes them as hard constraints will lead the generation to be biased towards a narrow distribution of objects and overfit to the training data. It causes limited generalization to unseen images. To alleviate this issue, we propose to train our model in a classifier-free manner by dropping $c$ and $G$ randomly at $50\%$ rate during training. In this way, the model is forced to generate objects that are visually consistent with the input image without being strictly constrained by them. Then $c$ can be regarded as an optional condition and $G$ as a condition secondary to the image input. At the same time, we also apply classifier-free training to $\mathcal{I}$ at $10\%$ dropping rate so that the image guidance can be further enhanced at inference by $\epsilon_{\phi}(\mathbf{x}_{t};\mathcal{I},G)=(1+\omega)\epsilon_{\phi}(\mathbf{x}_{t};t,\mathcal{I},G)-\omega\epsilon_{\phi}(\mathbf{x}_{t};t,\varnothing,G)$ , where $\omega$ is a hyperparameter to balance the influence of the image and the graph, and $\varnothing$ denotes the absence of the image.
@@ 3.3.3 Training Objectives
We supervise our model with the loss $\mathcal{L}=\mathcal{L}_{\epsilon}+\lambda\mathcal{L}_{\text{fg}}$, where $\mathcal{L}_{\epsilon}$ is noise residual loss and $\mathcal{L}_{\text{fg}}$ is object foreground loss. Our model follows the training procedure of the denoising diffusion probabilistic model (Ho et al., 2020), and $\mathcal{L}_{\epsilon}$ is the standard diffusion loss, which computes the mean squared error between the output $\epsilon_{t}$ and the ground truth residual noise $\hat{\epsilon}_{t}$ that is added to the input $\mathbf{x}_{t-1}$ in the timestep $t$ in the diffusion process. The loss $\mathcal{L}_{\text{fg}}$ is defined on the cross attention map $\mathcal{A}_{\mathbf{p}_{i}}^{l}$ for each part $\mathbf{p}_{i}$ at layer $l$, which records the attention scores from the part to all the image patches. $\mathcal{L}_{\text{fg}}$ is designed to encourage the part attention to focus on the foreground patches of the object across all parts and all layers. We define the loss as $\mathcal{L}_{\text{fg}}=\frac{1}{L}\sum_{l=1}^{L}{\sum_{i=1}^{N}{(1-\mathcal{A}_{\mathbf{p}_{i}}^{l}\circ\mathcal{M}_{\text{fg}}+\mathcal{A}_{\mathbf{p}_{i}}^{l}\circ\neg\mathcal{M}_{\text{fg}})}}$, where $\mathcal{M}_{fg}$ is the foreground mask of the object projected onto the patches, and $\circ$ denotes the element-wise product. We use $\lambda=0.1$ to balance the two losses.
@@ 4.1 Dataset
We collect data from PartNet-Mobility dataset (Xiang et al., 2020) to train our model across 7 categories (i.e., Storage, Table, Refrigerator, Dishwasher, Oven, Washer, and Microwave). We render 20 images per object from random viewpoints, sampled within a $90^{\circ}$ horizontal range and a $90^{\circ}$ upwards vertical range relative to the object’s front. With several augmentation strategies applied, we end up with 3,063 objects paired with 20 images rendered from random views for training, and additional 77 objects paired with 2 random views for testing. We also use 135 objects from the ACD dataset (Iliash et al., 2024) for additional evaluation in the zero-shot manner to test the generalization capability. In total, we have 55K training samples and 424 test samples in the experiments. Please see the appendix for more details on the data augmentation.
@@ 4.2 Baselines
We compare our method with two baselines: URDFormer (Chen et al., 2024b) as the state-of-the-art that works in the same setting and NAP (Lei et al., 2023) as representative prior work that unconditionally generates articulated objects. For URDFormer, we take their pretrained model for part detection on the input image and re-implement their object-level module to finetune on our training data for articulated object prediction. For NAP, we re-implement their method by plugging our image cross-attention module to attend with their node attributes after each of their graph attention layer to enable NAP to condition on an image. This variant of NAP is denoted by NAP-ICA. We also replace the shape latent originally used in NAP with part semantic label for sharing our retrieval algorithm, which is a lightweight version of NAP that has been shown to perform better in a conditional setting by Liu et al. (2024c). For comparison, all the methods are trained with or finetuned on the same training set and evaluated on the same test set. The final geometry of objects are from retrieval across all baseline methods. Please see the appendix for more implementation details.
@@ 4.3 Metrics
To evaluate the reconstruction accuracy in terms of shape structure and kinematics, we adopt 3 sets of evaluation metrics to measure the similarity between the output and the ground-truth object that correspond to the given image. For each metric set, the object similarity is measured both in the resting state (RS-) and across the articulated states (AS-) uniformly sampling from resting to the end state. The overall object scale is normalized to the one of the ground-truth objects to account for scale ambiguity across methods. Each metric is reported by averaging over the pairwise distance per part per state, and the part pairing is determined by the Hungarian algorithm.
@@ 4.4 Results
Reconstruction quality. We show the quantitative comparison on PartNet-Mobility dataset in Table 1 and qualitative results in Figure 4. In terms of the topological correctness of the predicted graph, our GPT-4o module achieves the highest accuracy. The graph prediction of URDFormer is less accurate and we observe that it is largely bottle-necked by their part detection module which often misses parts. NAP-ICA predicts more accurate graphs when node connectivity is not enforced as a constraint, which is consistent with the observation in CAGE that NAP struggles to adhere to graph conditioning. In terms of the geometric consistency with the input image, we observe that our method and NAP-ICA handle more variational input views and have better part-level understanding from the image than URDFormer, which leads to higher reconstruction quality by referring to the “RS-" metrics and qualitative results in Figure 4. This high performance is largely attributed to the ICA module we proposed that enables us and NAP to efficiently distill part knowledge from the rich DINO features. Our method further outperforms NAP-ICA, which is benefiting from the efficient graph guidance to better coordinate parts, especially when the number of parts gets larger. In terms of the plausibility of the part articulation, our method can generate more kinematically plausible objects with less part collisions by referring to the “AS-" and AOR metrics. It is important to note that the ’AS-’ metrics cannot fully capture motion plausibility, as there are cases where the ground-truth motion is not unique. From Figure 4, we observe that URDFormer exhibits inaccurate joint axis of doors due to its reliance on a predicted label that conflates semantic and joint categorizations, thereby impairing joint parameter estimation in their post-processing step. NAP-ICA tends to produce more collisions and less plausible motions which is also reflected in the user study below.
@@
Generalization to unseen dataset. To further evaluate the generalization capability, we show the comparison results on ACD dataset in a zero-shot manner. The ACD dataset comprises objects with greater structural complexity, geometric diversity, and appearance variances compared to the PartNet-Mobility dataset. Figure 5 shows examples from the ACD dataset, where our method generalize better to objects with more challenging textures and more complex part structures that are out-of-distribution. As shown in Table 2, we outperform other baselines overall by a large margin.
@@
Comparison with multi-view reconstruction. We compare with Real2Code (Mandi et al., 2024), a state-of-the-art method for articulated object reconstruction from multi-views. As their code is not fully open-sourced, we follow the evaluation protocol shown in their paper. Specifically, Real2Code reconstructs an object from 12 views observing the object in an opening state, which reveals much more geometric and kinematic information than our setting that only observes the object in a resting state from single view. We retrain our model on their data split with our augmented data and evaluate on their test set to ensure a fair comparison. From Table 3, we observe that our method achieves a competitive results on simple objects with less parts and can even outperform Real2Code on objects with more parts in complex structures. Also, our score gap between CD on the whole and its part on average is smaller than Real2Code, indicating that our method reconstructs better part geometry.
Ablation studies. We ablate the design components of our method as shown in Table 4, where the average of 5 samples per input given GT part connectivity graph is reported on the ACD dataset. On top of our base model, the ICA module plays the main role in enabling image conditioning by distilling DINO features to part attributes via cross-attention. Its effectiveness is also demonstrated by adding this module to NAP and achieving performance close to our method. Removing object category (cCF) as a hard constraint allows the model to share knowledge across categories, which enhances the efficiency of training data utilization and leads to improved generalization. $\mathcal{L}_{\text{fg}}$ yields a modest improvement in performance by concentrating attention on the object region. The classifier-free training for the image and graph guidance enhances the model to the best performance by modulating the distribution farther to respect the input image and generalize better to unseen data. We also ablate on how the augmented data affects the performance in the supplement, where the augmentation strategies are shown to be effective in improving the generalization capability.
@@
Failure cases and limitations. Some failure cases are shown in Figure 5 in the last two rows. We observe three failure modes in our method: 1) incorrect graph predictions from input images with challenging textures; 2) less accurate part geometry, particularly for small parts (e.g., knobs in the second last example); and 3) inaccurate part arrangement, often due to missing components in the predicted graph or when the object’s structure is highly cluttered and complex (see the last example). These failures are due to the low resolution of the input image that provides limited part details for our model to infer accurate part geometry and arrangement, especially for complex objects with large number of parts compacted in a small number of patch features. The unbalanced training on the data with less complex structure and simple textures also hinder the generalization ability of our model. Further augmentation to balance the object complexity and increase texture diversity will improve generalization to real settings for future work.

### v3

# Singapo: Single Image Controlled Generation of Articualted Parts in Objects

@@ 1 Introduction
Figure 1: Our work proposes to generate 3D articulated objects from a single image observing the object in the resting state from a random view. Left: We design a pipeline to synthesize articulated assets progressively from coarse to fine details in a modular way. Right: Our method generates objects with varying part geometry and motion to account for the ambiguity in the input image.
@@ 2 Related Work
Conditional generation on 3D structures. Enabling control over the generation process is crucial for practical applications. This becomes more challenging for structured data generation, due to the need for fine-grained control and parsing the conditioning into structured components. Part-based object generation has been studied in settings driven by a partial shape to complete the object part-by-part or guided by an image (Wu et al., 2020; Niu et al., 2018). These methods only work with specific categories of objects that share similar structures. The recent advances in large-language models and generative models promote conditional generation from 2D (Yang et al., 2021; Rombach et al., 2022; Zhang et al., 2023a; Blattmann et al., 2023) to 3D with structures (Liu et al., 2024a). Many works on 3D scene generation have explored the control of object arrangements via text for room and object specification (Zhang et al., 2023b; Lin & Mu, 2024; Aguina-Kang et al., 2024; Tam et al., 2024), 2D/3D layout for spatial constraint (Bahmani et al., 2023; Eldesokey & Wonka, 2024; Feng et al., 2024; Po & Wetzstein, 2024; Maillard et al., 2024; Yang et al., 2024), and scene graphs as spatial or semantic priors (Lin & Mu, 2024; Zhai et al., 2024; Gao et al., 2024). Chen et al. (2024a) generates objects from an image by decomposing the subscene into objects to generate separately. Alam & Ahmed (2024) propose to generate parametric CAD command sequences conditioned on an image. Our work follows the trend of conditional generation on structured data. To the best of our knowledge, we are the first to explore articulated object generation from a single image.
@@ 3 Methods
Figure 2: Our method takes an object image as input and generates attributes of articulated parts, which are used to assemble the object via part mesh retrieval. We design a diffusion-based model for part generation, which is guided by a part connectivity graph and DINOv2 patch features of the input image. Our denoising network is built on layers of attention blocks. The graph constraint is injected into the graph relation module by converting to an adjacency matrix as the attention mask. The image features act as the keys and values in the cross attention to condition the part arrangement.
@@ 3.2 Data Representation
We assume objects to be normalized inside a cube of size $2$ centered at the origin and all the parameters are defined in the world coordinate system. We describe each articulated part with multiple attributes $\mathbf{p}_{i}=\{\mathbf{b}_{i},l_{i},t_{i},\mathbf{a}_{i},\mathbf{r}_{i}\}$ at abstract level, including the part bounding box $\mathbf{b}_{i}\in\mathbb{R}^{6}$, semantic label $l_{i}$, articulation type $t_{i}$, location and direction of joint axis $\mathbf{a}_{i}\in\mathbb{R}^{6}$, and motion range ${\color[rgb]{0,0,0}\mathbf{r}_{i}}\in\mathbb{R}^{2}$. The articulation type considered in this work includes fixed, revolute, prismatic, continuous and screw joints. The semantic label includes base, door, drawer, handle, knob, tray. The semantic label will be used to retrieve the part geometry from the dataset. Spatially, the parts are organized into a connectivity graph, where each node represents a part, and the edge direction indicates the parent-child relationship. This spatial relationship also defines the kinematic constraints between parts. During training, the graph is represented as an adjacency matrix $A\in\{0,1\}^{N\times N}$, where $A_{ij}=1$ if part $i$ is in connection with part $j$. The matrix $A$ will serve as an attention mask to guide the part generation process. To handle a different number of parts in the denoising diffusion process, we pad the parts to a maximum number $32$ with all part attributes represented in $6$ dimensions.
@@ 3.3.2 Model architecture
Classifier-free training for conditioning. Our denoising network takes image $\mathcal{I}$ as the main control and graph $G$ as auxiliary guidance. Both of them are critical as the image describes the spatial arrangement of parts and the graph determines the ordering of the parts to be generated so that the parts are organized correctly when articulating. Inherent from the base model, our network can also take object category $c$ as an additional condition, which is fused with the timestep embeddings via an adaptive norm layer (Xu et al., 2019) prior to attentions in each block. We empirically find that the original design of CAGE that takes them as hard constraints will lead the generation to be biased towards a narrow distribution of objects and overfit to the training data. It causes limited generalization to unseen images. To alleviate this issue, we propose to train our model in a classifier-free manner by dropping $c$ and $G$ randomly at a $50\%$ rate during training. In this way, the model is forced to generate objects that are visually consistent with the input image without being strictly constrained by them. Then $c$ can be regarded as an optional condition and $G$ as a condition secondary to the image input. At the same time, we also apply classifier-free training to $\mathcal{I}$ at $10\%$ dropping rate so that the image guidance can be further enhanced at inference by $\epsilon_{\phi}(\mathbf{x}_{t};\mathcal{I},G)=(1+\omega)\epsilon_{\phi}(\mathbf{x}_{t};t,\mathcal{I},G)-\omega\epsilon_{\phi}(\mathbf{x}_{t};t,\varnothing,G)$ , where $\omega$ is a hyperparameter to balance the influence of the image and the graph, and $\varnothing$ denotes the absence of the image.
@@ 3.3.3 Training Objectives
We supervise our model with the loss $\mathcal{L}=\mathcal{L}_{\epsilon}+\lambda\mathcal{L}_{\text{fg}}$, where $\mathcal{L}_{\epsilon}$ is noise residual loss and $\mathcal{L}_{\text{fg}}$ is object foreground loss. Our model follows the training procedure of the denoising diffusion probabilistic model (Ho et al., 2020), and $\mathcal{L}_{\epsilon}$ is the standard diffusion loss, which computes the mean squared error between the output $\epsilon_{t}$ and the ground truth residual noise $\hat{\epsilon}_{t}$ that is added to the input $\mathbf{x}_{t-1}$ in the timestep $t$ in the diffusion process. $\mathcal{L}_{\text{fg}}$ is an auxiliary loss that helps focus the attention on the foreground patches of the object region for all parts and all layers. The loss $\mathcal{L}_{\text{fg}}$ supervises the cross attention map $\mathcal{A}_{\mathbf{p}_{i}}^{l}$ for part $\mathbf{p}_{i}$ at layer $l$, which records the attention scores from the part to all the image patches. Formally, we define the loss as $\mathcal{L}_{\text{fg}}=\frac{1}{L}\sum_{l=1}^{L}{\sum_{i=1}^{N}{(1-\mathcal{A}_{\mathbf{p}_{i}}^{l}\circ\mathcal{M}_{\text{fg}}+\mathcal{A}_{\mathbf{p}_{i}}^{l}\circ\neg\mathcal{M}_{\text{fg}})}}$, where $\mathcal{M}_{fg}$ is the foreground mask of the object projected onto the patches, and $\circ$ denotes the element-wise product. We use $\lambda=0.1$ to balance the two losses.
@@ 4.1 Dataset
We collect data from PartNet-Mobility dataset (Xiang et al., 2020) to train our model across 7 categories (i.e., Storage, Table, Refrigerator, Dishwasher, Oven, Washer, and Microwave). We render 20 images per object from random viewpoints, sampled within a $90^{\circ}$ horizontal range and a $90^{\circ}$ upwards vertical range relative to the object’s front. With several augmentation strategies applied, we end up with 3,063 objects paired with 20 images rendered at resting state for training, and additional 77 objects paired with 2 random views for testing. We also use 135 objects from the ACD dataset (Iliash et al., 2024) for additional evaluation in the zero-shot manner to test the generalization capability. In total, we have 55K training samples and 424 test samples in the experiments. Please see the appendix for more details on data augmentation.
@@ 4.2 Baselines
We compare our method with two baselines: URDFormer (Chen et al., 2024b) as the state-of-the-art that works in the same setting and NAP (Lei et al., 2023) as representative prior work that unconditionally generates articulated objects. For URDFormer, we take their pretrained model for part detection on the input image and re-implement their object-level module to finetune on our training data for articulated object prediction. For NAP, we re-implement their method by plugging our image cross-attention module to attend with their node attributes after each of their graph attention layer to enable NAP to condition on an image. This variant of NAP is denoted by NAP-ICA. We also replace the shape latent originally used in NAP with part semantic label for sharing our retrieval algorithm, which is a lightweight version of NAP that has been shown to perform better in a conditional setting by Liu et al. (2024c). For comparison, all the methods are trained with or finetuned on the same training set and evaluated on the same test set. The final geometry of objects is from retrieval across all baseline methods. Please see the appendix for more implementation details.
@@ 4.3 Metrics
To evaluate the reconstruction accuracy in terms of shape structure and kinematics, we adopt 3 sets of evaluation metrics to measure the similarity between the output and the ground-truth object that corresponds to the given image. For each metric set, the object similarity is measured both in the resting state (RS-) and across the articulated states (AS-) uniformly sampling from resting to the end state. The overall object scale is normalized to the one of the ground-truth objects to account for scale ambiguity across methods. Each metric is reported by averaging over the pairwise distance per part per state, and the part pairing is determined by the Hungarian algorithm.
@@ 4.4 Results
Reconstruction quality. We show the quantitative comparison on PartNet-Mobility dataset in Table 1 and qualitative results in Figure 4. In terms of the topological correctness of the predicted graph, our GPT-4o module achieves the highest accuracy. The graph prediction of URDFormer is less accurate, and we observe that it is largely bottlenecked by their part detection module which often misses parts. NAP-ICA predicts more accurate graphs when node connectivity is not enforced as a constraint, which is consistent with the observation in CAGE that NAP struggles to adhere to graph conditioning. In terms of the geometric consistency with the input image, we observe that our method and NAP-ICA handle more variational input views and have better part-level understanding from the image than URDFormer, which leads to higher reconstruction quality by referring to the “RS-" metrics and qualitative results in Figure 4. This high performance is largely attributed to the ICA module we proposed that enables us and NAP to efficiently distill part knowledge from the rich DINO features. Our method further outperforms NAP-ICA, which is benefiting from the efficient graph guidance to better coordinate parts, especially when the number of parts gets larger. In terms of the plausibility of the part articulation, our method can generate more kinematically plausible objects with less part collisions by referring to the “AS-" and AOR metrics. It is important to note that the ’AS-’ metrics cannot fully capture motion plausibility, as there are cases where the ground-truth motion is not unique. From Figure 4, we observe that URDFormer exhibits inaccurate joint axis of doors due to its reliance on a predicted label that conflates semantic and joint categorizations, thereby impairing joint parameter estimation in their post-processing step. NAP-ICA tends to produce more collisions and less plausible motions, which is also reflected in the user study below.
@@
Generalization to unseen dataset. To further evaluate the generalization capability, we show the comparison results on the ACD dataset in a zero-shot manner. The ACD dataset comprises objects with greater structural complexity, geometric diversity, and appearance variances compared to the PartNet-Mobility dataset. Figure 5 shows examples from the ACD dataset, where our method generalize better to objects with more challenging textures and more complex part structures that are out-of-distribution. As shown in Table 2, we outperform other baselines overall by a large margin.
@@
Comparison with multi-view reconstruction. We compare with Real2Code (Mandi et al., 2024), a state-of-the-art method for articulated object reconstruction from multi-views. As their code is not fully open-sourced, we follow the evaluation protocol shown in their paper. Specifically, Real2Code reconstructs an object from 12 views observing the object in an opening state, which reveals much more geometric and kinematic information than our setting that only observes the object in a resting state from a single view. We retrain our model on their data split with our augmented data and evaluate on their test set to ensure a fair comparison. From Table 3, we observe that our method achieves competitive results on simple objects with less parts and can even outperform Real2Code on objects with more parts in complex structures. Also, our score gap between CD on the whole and its part on average is smaller than Real2Code, indicating that our method reconstructs better part geometry.
Ablation studies. We ablate the design components of our method as shown in Table 4, where the average of 5 samples per input given GT part connectivity graph is reported on the ACD dataset. On top of our base model, the ICA module plays the main role in enabling image conditioning by distilling DINOv2 features to part attributes via cross-attention. Its effectiveness is also demonstrated by adding this module to NAP and achieving performance close to our method. Removing object category (cCF) as a hard constraint allows the model to share knowledge across categories, which enhances the efficiency of training data utilization and leads to improved generalization. $\mathcal{L}_{\text{fg}}$ yields a modest improvement in performance by concentrating attention on the object region. The classifier-free training for the image and graph guidance enhances the model to the best performance by modulating the distribution further to respect the input image and generalize better to unseen data. We also ablate on how the augmented data affects the performance in the supplement, where the augmentation strategies are shown to be effective in improving the generalization capability.
@@
Failure cases and limitations. Some failure cases are shown in Figure 5 in the last two rows. We observe three failure modes in our method: 1) incorrect graph predictions from input images with challenging textures; 2) less accurate part geometry, particularly for small parts (e.g., knobs in the second last example); and 3) inaccurate part arrangement, often due to missing components in the predicted graph or when the object’s structure is highly cluttered and complex (see the last example). These failures are due to the low resolution of the input image that provides limited part details for our model to infer accurate part geometry and arrangement, especially for complex objects with a large number of parts compacted in a small number of patch features. The unbalanced training on the data with less complex structure and simple textures also hinder the generalization ability of our model. Further augmentation to balance the object complexity and increase texture diversity will improve generalization to real settings for future work.